# A proposal for an incident regime that tracks and counters threats to national security posed by AI systems

### v1

@@ Abstract
Recent progress in AI capabilities has heightened concerns that AI systems could pose a threat to national security, for example, by making it easier for malicious actors to perform cyberattacks on critical national infrastructure or through the loss of control of autonomous AI systems. In parallel, several national legislatures, including the US, have proposed nascent ‘AI incident reporting regimes’ to identify and counter similar threats. In this paper, we consolidate these two trends and present a proposal for a legally mandated post-deployment AI incident regime that aims to counter potential national security threats from AI systems. We start the paper by introducing the concept of ‘security-critical’ to describe sectors that pose extreme risks to national security, before arguing that civilian nuclear power, aviation, and life sciences dual-use research of concern, and frontier AI development, should all be considered ‘security-critical’ sectors. We then present in detail our AI incident proposal, justifying each component of the proposal in by demonstrating its similarity to US domestic incident regimes in other ‘security-critical’ sectors. Finally, we sketch a hypothetical scenario where our proposed AI incident regime deals with an AI cyber incident. Our proposed AI incident regime is split into three phases: in the first phase, we propose a novel operationalisation of what counts as an ‘AI incident’: we suggest that AI providers must create ‘national security cases’ for their frontier AI systems before they are deployed, and that AI providers must report as an incident any event which weakens a claim made in a ‘national security case’ for the AI system. In the second phase, we suggest that, once an AI provider has discovered an incident has occurred, the provider should notify a government agency at most 24 hours later, so that the government agency can coordinate an emergency containment response if appropriate. In the third and final phase, we suggest that, in order to counter future threats to national security, the government agency should be granted the ‘intelligence-gathering authority’ to obtain sufficient information to investigate the incident, and a ‘security-strengthening authority’ to demand all relevant AI providers to amend their security and safety procedures. In using incident reports to spur improvements to AI providers’ security and safety procedures, our AI incident regime proposal is grounded in established practice from incident regimes in other security-critical sectors and is a timely intervention given ongoing policy interest in the potential national security threats posed by AI systems.
@@ 5 Conclusion
We have put forward, and justified, a proposal for an AI incident regime that can help counter potential post-public deployment national security threats from AI systems. We started the paper by coining the term ‘security-critical’ to describe sectors that pose extreme risks to national security, and argue that the term ‘security-critical’ describes civilian nuclear power, aviation, and life sciences dual-use research of concern, and frontier AI development. Then, the main part of the paper consisted of detailing our proposal for an AI incident regime. We justified each component of our AI incident proposal by demonstrating that it mirrors the US incident regimes in other security-critical sectors. Our proposal had three phases; first, prior to public deployment, AI providers must create national security cases for their frontier AI systems before they are deployed, after which they must report as an incident any event which weakens a claim made in a ‘national security case’ for the AI system. Secondly, the ‘Rapid Response Phase’: providers must must notify a government agency no later than 24 hour after they discover that an incident has occurred, so that the government agency can coordinate an emergency containment response if appropriate. And finally, the ‘Hardening Defenses Phase’ involves countering future threats to national security. Here, recommend that the government agency have two authorities – an ‘intelligence-gathering authority’ which gives the government agency sufficient access to investigate the incident, and a ‘security-strengthening authority’ which allows the government agency to direct all relevant AI providers to amend their security and safety procedures. Against a backdrop where AI systems’ ability to pose national security threats has rapidly progressed, and may continue to increase, the three phases of our proposal are grounded in common practice in other security-critical sectors, offering an intervention that can help to track and counter the national security threats that may be posed by AI systems.

### v2

# Countering threats to national security posed by AI systems through an incident regime

@@ Abstract
Recent progress in AI capabilities has heightened concerns that AI systems could pose a threat to national security, for example, by making it easier for malicious actors to perform cyberattacks on critical national infrastructure, or through loss of control of autonomous AI systems. In parallel, federal legislators in the US have proposed nascent ’AI incident regimes’ to identify and counter similar threats. In this paper, we consolidate these two trends and present a timely proposal for a legally mandated post-deployment AI incident regime that aims to counter potential national security threats from AI systems. We start the paper by introducing the concept of ’security-critical’ to describe doctors that pose extreme risks to national security, before arguing that ’security-critical’ describes civilian nuclear power, aviation, life science dual-use research of concern, and frontier AI development. We then present in detail our AI incident regime proposal, justifying each component of the proposal by demonstrating its similarity to US domestic incident regimes in other ’security-critical’ sectors. Finally, we sketch a hypothetical scenario where our proposed AI incident regime deals with an AI cyber incident. Our proposed AI incident regime is split into three phases. The first phase revolves around a novel operationalization of what counts as an ’AI incident’ and we suggest that AI providers must create a ’national security case’ before deploying a frontier AI system. The second and third phases spell out that AI providers should notify a government agency about incidents, and that the government agency should be involved in amending AI providers’ security and safety procedures, in order to counter future threats to national security.
@@ 5 Conclusion
We have put forward, and justified, a proposal for an AI incident regime that can help counter potential post-public deployment national security threats from AI systems. We started the paper by coining the term ‘security-critical’ to describe sectors that pose extreme risks to national security, and argued that the term ‘security-critical’ describes civilian nuclear power, aviation, and life sciences dual-use research of concern, and frontier AI development. Then, the main part of the paper consisted of detailing our proposal for an AI incident regime. We justified each component of our AI incident proposal by demonstrating that it mirrors the US incident regimes in other security-critical sectors. Our proposal had three phases; first, prior to public deployment, AI providers must create national security cases for their frontier AI systems before they are deployed, after which they must report as an incident any event which weakens a claim made in a ‘national security case’ for the AI system. Secondly, the ‘Rapid Response Phase’: providers must must notify a government agency no later than 24 hour after they discover that an incident has occurred, so that the government agency can coordinate an emergency containment response if appropriate. And finally, the ‘Hardening Defenses Phase’ involves countering future threats to national security. Here, recommend that the government agency have two authorities – an ‘intelligence-gathering authority’ which gives the government agency sufficient access to investigate the incident, and a ‘security-strengthening authority’ which allows the government agency to direct all relevant AI providers to amend their security and safety procedures. Against a backdrop where AI systems’ ability to pose national security threats has rapidly progressed, and may continue to increase, the three phases of our proposal are grounded in common practice in other security-critical sectors, offering an intervention that can help to track and counter the national security threats that may be posed by AI systems.